# MeAJOR Corpus: A Multi-Source Dataset for Phishing Email Detection

### v1

@@ Abstract
Index Terms:
phishing, dataset, email, artificial intelligence, machine learning, deep learning
@@ I Introduction
Phishing remains one of the most prevalent and damaging forms of cybercrime, counting with over 3.76 million cyberattacks in 2024, causing wide financial losses globally and contributing to hundreds of thousands of compromised accounts (mostly email) in 2024 [1, 2, 3, 4]. It is also responsible for 15% of attack vectors in data breaches, costing an average of \$4.88 million per breach [5]. Email is one of the many media through which these attacks are performed, as it remains a central channel for personal and professional communication, and with the ever-evolving attack landscape, automated detection and mitigation strategies have become essential to safeguarding users and infrastructures. Machine Learning (ML) approaches promise scalable and adaptive defence mechanisms against evolving phishing campaigns, however, their effectiveness is highly dependent on the quality and diversity of the training data.
Despite the availability of several open-source phishing datasets, recent literature highlights some limitations in their usability for robust model development [6]. These issues include extreme class imbalance, limited feature engineering, narrow coverage of phishing tactics, outdated samples, and inconsistent preprocessing practices [6, 7, 8]. Furthermore, many studies rely on private or small-scale datasets, which limits generalisability and reproducibility [9]. These challenges restrain the development of efficient and realistic ML models, as they may lead to overfitting or inflated performance metrics on small or biased test sets.
@@ II-A Data Sources
There are several publicly available email spam and phishing datasets, some of which have been used for ML-based email phishing detection [10]. Spam datasets are commonly used in phishing detection studies [11, 12] because phishing is a type of spam and these emails share many characteristics, such as being unsolicited, unwanted, deceptive, and attempting to solicit an action [13]. Table I provides an overview of these datasets, including their content type, label availability, and sample sizes.
@@
The corpora available for phishing detection vary significantly in origin and content. They range from large, unlabelled corporate archives like the Enron Corpus [14], often used as a source of benign emails, to specialised collections containing only malicious samples, such as the Nazario Phishing Corpus [15] and Nigerian Fraud letters [16]. Many were curated specifically for academic research and benchmarking, including the seminal Ling-Spam [17], the SpamAssassin corpus [18], and collections from standardised competitive events like the TREC datasets [19, 20, 21] and the CEAS-08 [22]. More recent efforts focus on developing labelled, phishing-specific datasets, for example by extracting and curating samples from larger collections, as demonstrated by initiatives like Email4S [23].
Existing phishing email datasets face significant challenges that restrict the development of robust ML-based detection systems. Recent reviews highlight that many corpora suffer from data imbalance, outdated structures, and limited feature engineering. There is often a trade-off between scope and diversity; classic spam corpora like the SpamAssassin dataset offer high source diversity, whereas dedicated phishing collections, such as the Nazario Phishing Corpus, tend to be smaller and narrower in scope, lacking coverage of varied phishing tactics. These differences can heavily impact study results [7].
Furthermore, much of the current research relies on small public or inaccessible private datasets, which affects generalisability. Ethical concerns, such as the disclosure of sensitive or personally identifiable information, also make more difficult to acquire new datasets. Although some studies attempt to mitigate these issues by merging multiple corpora, these efforts often remain limited in scale, combining only two or three sources [24]. Addressing these quality issues is crucial for improving the reliability and generalisability of phishing detection research [8, 6].
@@ II-B Phishing Email Features
In the phishing email domain, researchers commonly extract features from six principal sources: email body text, embedded URLs, attachments, message headers, HTML structure, and external domain reputation, to expose both technical anomalies (e.g. mismatched “From” headers or IP-based links) and social-engineering cues (e.g. urgency words or deceptive phrasing) [25, 26, 27, 28].
@@
Email Body Text Features contain free-form writing that can reveal many social-engineering tactics and tricks used by attackers to exploit the human factor of a system [29, 30]. These cues are evident not only in the direct content of the email but also in its lexicality, syntax, and semantics. Some approaches used to extract these features from text include vectorisation methods such as Bag of Words or TF-IDF [31], counts of trigger words like “verify,” “password,” and “urgent” [32], readability scores such as Flesch–Kincaid and Gunning Fog [33], and sentiment analysis to detect emotional manipulation [34]. Together, these linguistic characteristics help identify deceptive emails crafted to manipulate recipients.
@@
Embedded URL Features are crucial in identifying phishing emails, as attackers often use deceptive links to lure victims. These features typically include various metrics related to the URL’s lexical structure, such as the total length of the URL, the number of subdomains it contains, the presence of IP-literal hostnames (e.g. https://192.168.1.1/home), the directory depth (measured by counting the segments separated by “/”), token counts derived by splitting the URL on characters like “.”, “-”, and “_”, digit‑to‑letter ratios and the overall character entropy. Beyond these lexical characteristics, other suspicious indicators include the use of dubious top-level domains, the presence of UTF-encoded characters, and the employment of excessive URL-shortening services [35, 36, 37]. These combined URL features help in detecting potentially harmful links embedded in phishing emails.
@@
Attachment Features play a vital role in detecting malicious content within emails, as attachments can carry malicious payloads or impersonate legitimate forms. These features include the number and types of attached files (e.g. executables, Office documents containing macros, and compressed archives), file‐size anomalies, discrepancies between the file extension and the MIME type, embedded macros or scripts, and the count of embedded images. Additionally, statistical flags for unusually large or zero-byte attachments and signature-based checks on macro patterns are effective tools for identifying concealed malware or credential-harvesting forms [38, 39]. Together, these features help uncover malicious attachments that aim to compromise the recipient’s security.
@@
Message Header Features carry email’s transport and routing metadata, which attackers often spoof or manipulate to conceal their identity or mislead recipients. Notable examples include the number and sequence of the Received header fields, which records each server the email passed through and creates a chain with timestamps (abnormal counts or irregular sequences in these hops may indicate suspicious routing or email relaying through unexpected servers); discrepancies between the From header field, which shows the sender claimed email; the Return-Path field, which specifies the destination for bounce messages; and the envelope sender (the actual sender address used during SMTP transmission), which can reveal inconsistencies or attempts to mask the sender identity. Authentication results from protocols like SPF (Sender Policy Framework), DKIM (DomainKeys Identified Mail), and DMARC (Domain-based Message Authentication, Reporting & Conformance) also provide strong spoofing indicators by verifying if emails are actually from who they claim to be from. Additional suspicious indicators include anomalous domains in the Reply-To field (which specifies the address that replies will be directed to) or List-Unsubscribe field (used for newsletter opt-outs), as attackers might use these fields to redirect victims or avoid detection; and unusual timestamp discrepancies, such as impossible or inconsistent sending times across headers. By quantifying these signals, detection models can effectively learn to identify forged or illicit senders and flag potentially malicious emails [40].
@@
HTML Structure Features are essential for analysing emails that include HTML content, as attackers often embed hidden forms, scripts, or obfuscated elements to deceive recipients. These features involve counting and examining HTML tags commonly exploited in phishing attempts, such as <form> (which can collect user credentials), <iframe> (often used to load external content invisibly), <script> (which may execute malicious code), <img> (which can track user interaction or load deceptive visuals), and inline CSS (frequently used to hide elements from the recipient). Key structural indicators include the depth of tag nesting (which can suggest attempts to hide content deep within complex layouts), the ratio of hidden elements (e.g. those using display:none) to visible ones, mismatches between anchor text and the actual href targets (which can disguise malicious links), and the presence of JavaScript event handlers like onload or onclick (which can trigger actions when the user interacts with the email). By extracting and analysing these elements, it becomes possible to uncover the client-side techniques used by phishing kits to bypass basic text-based detection methods [41, 42].
@@
Reputation Features extend beyond the immediate content of an email to evaluate the trustworthiness of the attacker’s domains. Important reputation indicators include the domain’s registration age, time remaining until domain expiration, domain name registrar and domain nameserver metadata (which can reveal connections to known malicious actors or unreliable service providers), DNS record time-to-live values (which indicate the refresh rate of domain information), and the geographic location of the hosting IP address (adds further context, as certain regions may be associated with phishing or cybercrime). Additionally, it is crucial to check whether the host IP or Autonomous System Number (ASN) is on known Blacklists. Domains with short lifespans (newly registered or rapidly changing), use obscure top-level domains, or rely on hosting providers with lax abuse policies tend to have a strong association with phishing campaigns. These reputation features provide valuable signals for identifying potentially malicious domains used in attacks [43, 44].
@@ II-C Phishing Detection Techniques
There are many techniques used for phishing email detection. Some of the most traditional ones include reputation-based methods, which involve assessing the trustworthiness of a sender or website based on historical data [45]; heuristic-based methods, which rely on flexible, experience-informed criteria [46]; and rule-based methods, which classify content according to predefined rules [12, 47]. Despite having consistent detection rates for known threats, these methods lack zero-day attack detection and are hard to maintain, which led researchers to search for novel approaches, often based on Natural Language Processing (NLP), ML, and DL.
@@
Omotehinwa et al. [48] evaluated ensemble ML models for spam email detection, emphasising hyperparameter optimisation. They trained RF and Extreme Gradient Boosting (XGB) models on textual features extracted from the Enron-Spam corpus. XGB performed best with an F1 of 98.16%, demonstrating the effectiveness of ensemble learning in identifying email threats.
@@
Naswir et al. [49] investigated how effectively lexical URL features can detect phishing emails. They extracted features like IP addresses in URLs, special characters (e.g. ’@’, ’-’, and ‘%’), URL entropy, and length, alongside basic email metadata, from the Nazario Phishing Corpus and the legitimate Enron Corpus. By optimising a Support Vector Machine (SVM) with the Cuckoo Search algorithm, their method achieved 91% accuracy, demonstrating the strong predictive power of URL-based features for phishing detection.
@@
Khalid et al. [11] introduced LogiTriBlend, a stacked ensemble model for phishing email detection, combining SVM, Logistic Regression, RF, and XGB. They used a small, imbalanced ”Phishing Email Dataset” (4591 emails) and addressed this imbalance with SMOTE data augmentation. It comprises Enron Corpus, SpamAssassin, Nazario Phishing Corpus and other sources. The models were trained on textual features, vectorised using TF-IDF, Word2Vec, and Doc2Vec. The Doc2Vec-based features with LogiTriBlend achieved the best F1 (99.41%). However, the study’s generalisability might be limited due to the small size of the original dataset and its reliance solely on text features. Chanis et al. [50] also used ensemble methods for phishing detection, uniquely combining stylometric features (such as formatting and writing style) with content-based features. They trained separate classifiers for each feature type, then integrated them using a stacked ensemble. This approach captured both the email’s meaning and writing style, achieving a 98.43% F1 on a private dataset, outperforming content-only models by around 2.2%.
@@
Atawneh and Aljehani [51] compared various Deep Learning (DL) architectures (CNNs, RNNs, LSTMs) for phishing detection, using textual features from email bodies from the Nazario and SpamAssassin corpora. Their proposed BERT+LSTM hybrid model achieved a 99.55% F1, outperforming other standalone models. Kaushik et al. [52] proposed another hybrid DL model, combining LSTM and CNN layers, for phishing detection. Their model processes URL-based and textual features (transformed into 2D image-like matrices) from a balanced dataset, constructed by combining phishing URL data from PhishTank [53] and legitimate URL data from Alexa [54] with additional email datasets from various public repositories. It achieved an F1 of around 99%, demonstrating the efficacy of combining sequential and spatial feature learning.
@@
Despite most previously mentioned studies using text-based features, other information can also be extracted and used as input for the models. For example, Zhang et al. [55] developed a phishing detection framework that uses diverse features beyond just text. Their approach extracts text from embedded images via Optical Character Recognition (OCR), combines it with URL-based, web-based, and rule-based features. Using these varied inputs, they trained a two-stage Extreme Learning Machine (ELM) model on a dataset of over 22000 webpages collected from PhishTank, APWG, and legitimate sources such as Alexa’s top sites, achieving 99.04% accuracy.
@@
Zhang et al. [56] used 79 static email features (e.g., sender IP, links, attachments) and textual embeddings to train a RF classifier for phishing detection. Their model achieved 99.97% accuracy on a large, custom dataset (660985 emails), combining benign emails from Enron Corpus (517401 samples), Trec-07 (25220 samples), and SpamAssassin dataset (6952 samples); and malicious phishing samples from Nazario Phishing Corpus (9510 samples), Wooyun XSS dataset [57] (168 samples), Trec-07 (50199 samples), and additional 49136 samples generated by summarising and simulating novel types of malicious email attacks that exploit email protocol vulnerabilities. While this dataset is extensive, its heavy reliance on the Enron Corpus for benign emails might introduce bias.
@@
Altwaijry et al. [12] compared DL models for phishing email detection, using only email body and subject content from Nazario Phishing Corpus and SpamAssassin. They proposed a lightweight 1D-CNN architecture (1D-CNNPD), enhanced with recurrent layers like Bi-GRU. Their best model (1D-CNNPD + Bi-GRU) achieved a 99.66% F1, outperforming traditional ML models and comparable DL approaches such as THEMIS [58] and DeepAnti-PhishNet [59] and demonstrating the effectiveness of CNNs combined with recurrent mechanisms for phishing detection from text. Another recent approach by Koide et al. [60] introduced ChatSpamDetector, which uses Large Language Models (LLMs), specifically GPT-4, for phishing detection. They prompt GPT-4 directly with raw email text, allowing it to infer phishing characteristics contextually, rather than relying on engineered features. This zero-shot/few-shot approach achieved 99.70% accuracy on a private balanced dataset, showing LLMs’ ability to capture subtle phishing signals.

### v2

@@ I Introduction
Phishing remains one of the most prevalent and damaging forms of cybercrime, counting with over 3.76 million cyberattacks in 2024, causing wide financial losses globally and contributing to hundreds of thousands of compromised accounts (mostly email) in 2024 [apwg_q1, apwg_q2, apwg_q3, apwg_q4]. It is also responsible for 15% of attack vectors in data breaches, costing an average of \$4.88 million per breach [IBM_data_breach_report_2024]. Email is one of the many media through which these attacks are performed, as it remains a central channel for personal and professional communication, and with the ever-evolving attack landscape, automated detection and mitigation strategies have become essential to safeguarding users and infrastructures. Machine Learning (ML) approaches promise scalable and adaptive defence mechanisms against evolving phishing campaigns, however, their effectiveness is highly dependent on the quality and diversity of the training data.
Despite the availability of several open-source phishing datasets, recent literature highlights some limitations in their usability for robust model development [Mbadiwe2024]. These issues include extreme class imbalance, limited feature engineering, narrow coverage of phishing tactics, outdated samples, and inconsistent preprocessing practices [Mbadiwe2024, Alhuzali2025, AlSubaiey2024]. Furthermore, many studies rely on private or small-scale datasets, which limits generalisability and reproducibility [afonso2025rethinking]. These challenges restrain the development of efficient and realistic ML models, as they may lead to overfitting or inflated performance metrics on small or biased test sets.
@@ II-A Data Sources
There are several publicly available email spam and phishing datasets, some of which have been used for ML-based email phishing detection [Salloum2022]. Spam datasets are commonly used in phishing detection studies [Khalid2024, Altwaijry2024] because phishing is a type of spam and these emails share many characteristics, such as being unsolicited, unwanted, deceptive, and attempting to solicit an action [Juneja2014]. Table I provides an overview of these datasets, including their content type, label availability, and sample sizes.
@@
The corpora available for phishing detection vary significantly in origin and content. They range from large, unlabelled corporate archives like the Enron Corpus [enron_source], often used as a source of benign emails, to specialised collections containing only malicious samples, such as the Nazario Phishing Corpus [nazario_source] and Nigerian Fraud letters [nigerian_source]. Many were curated specifically for academic research and benchmarking, including the seminal Ling-Spam [ling_source], the SpamAssassin corpus [spamassassin_source], and collections from standardised competitive events like the TREC datasets [trec5source, trec6source, trec7source] and the CEAS-08 [ceas8source]. More recent efforts focus on developing labelled, phishing-specific datasets, for example by extracting and curating samples from larger collections, as demonstrated by initiatives like Email4S [email4s_source].
Existing phishing email datasets face significant challenges that restrict the development of robust ML-based detection systems. Recent reviews highlight that many corpora suffer from data imbalance, outdated structures, and limited feature engineering. There is often a trade-off between scope and diversity; classic spam corpora like the SpamAssassin dataset offer high source diversity, whereas dedicated phishing collections, such as the Nazario Phishing Corpus, tend to be smaller and narrower in scope, lacking coverage of varied phishing tactics. These differences can heavily impact study results [Alhuzali2025].
Furthermore, much of the current research relies on small public or inaccessible private datasets, which affects generalisability. Ethical concerns, such as the disclosure of sensitive or personally identifiable information, also make more difficult to acquire new datasets. Although some studies attempt to mitigate these issues by merging multiple corpora, these efforts often remain limited in scale, combining only two or three sources [Kyaw2024]. Addressing these quality issues is crucial for improving the reliability and generalisability of phishing detection research [AlSubaiey2024, Mbadiwe2024].
@@ II-B Phishing Email Features
In the phishing email domain, researchers commonly extract features from six principal sources: email body text, embedded URLs, attachments, message headers, HTML structure, and external domain reputation, to expose both technical anomalies (e.g. mismatched “From” headers or IP-based links) and social-engineering cues (e.g. urgency words or deceptive phrasing) [Lee2025, Gascon2018, Gualberto2020, Valecha2022].
@@
Email Body Text Features contain free-form writing that can reveal many social-engineering tactics and tricks used by attackers to exploit the human factor of a system [Desolda2021, Gallo2024]. These cues are evident not only in the direct content of the email but also in its lexicality, syntax, and semantics. Some approaches used to extract these features from text include vectorisation methods such as Bag of Words or TF-IDF [Tawil2024], counts of trigger words like “verify,” “password,” and “urgent” [Yasin2016], readability scores such as Flesch–Kincaid and Gunning Fog [Mieronkoski2020], and sentiment analysis to detect emotional manipulation [Xiang2025]. Together, these linguistic characteristics help identify deceptive emails crafted to manipulate recipients.
@@
Embedded URL Features are crucial in identifying phishing emails, as attackers often use deceptive links to lure victims. These features typically include various metrics related to the URL’s lexical structure, such as the total length of the URL, the number of subdomains it contains, the presence of IP-literal hostnames (e.g. https://192.168.1.1/home), the directory depth (measured by counting the segments separated by “/”), token counts derived by splitting the URL on characters like “.”, “-”, and “_”, digit‑to‑letter ratios and the overall character entropy. Beyond these lexical characteristics, other suspicious indicators include the use of dubious top-level domains, the presence of UTF-encoded characters, and the employment of excessive URL-shortening services [Joshi2019, Ahmed2022, Xie2022]. These combined URL features help in detecting potentially harmful links embedded in phishing emails.
@@
Attachment Features play a vital role in detecting malicious content within emails, as attachments can carry malicious payloads or impersonate legitimate forms. These features include the number and types of attached files (e.g. executables, Office documents containing macros, and compressed archives), file‐size anomalies, discrepancies between the file extension and the MIME type, embedded macros or scripts, and the count of embedded images. Additionally, statistical flags for unusually large or zero-byte attachments and signature-based checks on macro patterns are effective tools for identifying concealed malware or credential-harvesting forms [Chen2023, Rudd2018]. Together, these features help uncover malicious attachments that aim to compromise the recipient’s security.
@@
Message Header Features carry email’s transport and routing metadata, which attackers often spoof or manipulate to conceal their identity or mislead recipients. Notable examples include the number and sequence of the Received header fields, which records each server the email passed through and creates a chain with timestamps (abnormal counts or irregular sequences in these hops may indicate suspicious routing or email relaying through unexpected servers); discrepancies between the From header field, which shows the sender claimed email; the Return-Path field, which specifies the destination for bounce messages; and the envelope sender (the actual sender address used during SMTP transmission), which can reveal inconsistencies or attempts to mask the sender identity. Authentication results from protocols like SPF (Sender Policy Framework), DKIM (DomainKeys Identified Mail), and DMARC (Domain-based Message Authentication, Reporting & Conformance) also provide strong spoofing indicators by verifying if emails are actually from who they claim to be from. Additional suspicious indicators include anomalous domains in the Reply-To field (which specifies the address that replies will be directed to) or List-Unsubscribe field (used for newsletter opt-outs), as attackers might use these fields to redirect victims or avoid detection; and unusual timestamp discrepancies, such as impossible or inconsistent sending times across headers. By quantifying these signals, detection models can effectively learn to identify forged or illicit senders and flag potentially malicious emails [Kulkarni2020].
@@
HTML Structure Features are essential for analysing emails that include HTML content, as attackers often embed hidden forms, scripts, or obfuscated elements to deceive recipients. These features involve counting and examining HTML tags commonly exploited in phishing attempts, such as <form> (which can collect user credentials), <iframe> (often used to load external content invisibly), <script> (which may execute malicious code), <img> (which can track user interaction or load deceptive visuals), and inline CSS (frequently used to hide elements from the recipient). Key structural indicators include the depth of tag nesting (which can suggest attempts to hide content deep within complex layouts), the ratio of hidden elements (e.g. those using display:none) to visible ones, mismatches between anchor text and the actual href targets (which can disguise malicious links), and the presence of JavaScript event handlers like onload or onclick (which can trigger actions when the user interacts with the email). By extracting and analysing these elements, it becomes possible to uncover the client-side techniques used by phishing kits to bypass basic text-based detection methods [Yoon2024, Yoon2025].
@@
Reputation Features extend beyond the immediate content of an email to evaluate the trustworthiness of the attacker’s domains. Important reputation indicators include the domain’s registration age, time remaining until domain expiration, domain name registrar and domain nameserver metadata (which can reveal connections to known malicious actors or unreliable service providers), DNS record time-to-live values (which indicate the refresh rate of domain information), and the geographic location of the hosting IP address (adds further context, as certain regions may be associated with phishing or cybercrime). Additionally, it is crucial to check whether the host IP or Autonomous System Number (ASN) is on known Blacklists. Domains with short lifespans (newly registered or rapidly changing), use obscure top-level domains, or rely on hosting providers with lax abuse policies tend to have a strong association with phishing campaigns. These reputation features provide valuable signals for identifying potentially malicious domains used in attacks [Hao2016, Chiba2025].
@@ II-C Phishing Detection Techniques
There are many techniques used for phishing email detection. Some of the most traditional ones include reputation-based methods, which involve assessing the trustworthiness of a sender or website based on historical data [Kang2007]; heuristic-based methods, which rely on flexible, experience-informed criteria [Prakash2010]; and rule-based methods, which classify content according to predefined rules [Altwaijry2024, Doshi2023]. Despite having consistent detection rates for known threats, these methods lack zero-day attack detection and are hard to maintain, which led researchers to search for novel approaches, often based on Natural Language Processing (NLP), ML, and DL.
@@
Omotehinwa et al. [Omotehinwa2023] evaluated ensemble ML models for spam email detection, emphasising hyperparameter optimisation. They trained RF and Extreme Gradient Boosting (XGB) models on textual features extracted from the Enron-Spam corpus. XGB performed best with an F1 of 98.16%, demonstrating the effectiveness of ensemble learning in identifying email threats.
@@
Naswir et al. [Naswir2022] investigated how effectively lexical URL features can detect phishing emails. They extracted features like IP addresses in URLs, special characters (e.g. ’@’, ’-’, and ‘%’), URL entropy, and length, alongside basic email metadata, from the Nazario Phishing Corpus and the legitimate Enron Corpus. By optimising a Support Vector Machine (SVM) with the Cuckoo Search algorithm, their method achieved 91% accuracy, demonstrating the strong predictive power of URL-based features for phishing detection.
@@
Khalid et al. [Khalid2024] introduced LogiTriBlend, a stacked ensemble model for phishing email detection, combining SVM, Logistic Regression, RF, and XGB. They used a small, imbalanced ”Phishing Email Dataset” (4591 emails) and addressed this imbalance with SMOTE data augmentation. It comprises Enron Corpus, SpamAssassin, Nazario Phishing Corpus and other sources. The models were trained on textual features, vectorised using TF-IDF, Word2Vec, and Doc2Vec. The Doc2Vec-based features with LogiTriBlend achieved the best F1 (99.41%). However, the study’s generalisability might be limited due to the small size of the original dataset and its reliance solely on text features. Chanis et al. [Chanis2024] also used ensemble methods for phishing detection, uniquely combining stylometric features (such as formatting and writing style) with content-based features. They trained separate classifiers for each feature type, then integrated them using a stacked ensemble. This approach captured both the email’s meaning and writing style, achieving a 98.43% F1 on a private dataset, outperforming content-only models by around 2.2%.
@@
Atawneh and Aljehani [Atawneh2023] compared various Deep Learning (DL) architectures (CNNs, RNNs, LSTMs) for phishing detection, using textual features from email bodies from the Nazario and SpamAssassin corpora. Their proposed BERT+LSTM hybrid model achieved a 99.55% F1, outperforming other standalone models. Kaushik et al. [Kaushik2023] proposed another hybrid DL model, combining LSTM and CNN layers, for phishing detection. Their model processes URL-based and textual features (transformed into 2D image-like matrices) from a balanced dataset, constructed by combining phishing URL data from PhishTank [PhishTank] and legitimate URL data from Alexa [AlexaTopSites] with additional email datasets from various public repositories. It achieved an F1 of around 99%, demonstrating the efficacy of combining sequential and spatial feature learning.
@@
Despite most previously mentioned studies using text-based features, other information can also be extracted and used as input for the models. For example, Zhang et al. [Zhang2016] developed a phishing detection framework that uses diverse features beyond just text. Their approach extracts text from embedded images via Optical Character Recognition (OCR), combines it with URL-based, web-based, and rule-based features. Using these varied inputs, they trained a two-stage Extreme Learning Machine (ELM) model on a dataset of over 22000 webpages collected from PhishTank, APWG, and legitimate sources such as Alexa’s top sites, achieving 99.04% accuracy.
@@
Zhang et al. [Zhang2025] used 79 static email features (e.g., sender IP, links, attachments) and textual embeddings to train a RF classifier for phishing detection. Their model achieved 99.97% accuracy on a large, custom dataset (660985 emails), combining benign emails from Enron Corpus (517401 samples), Trec-07 (25220 samples), and SpamAssassin dataset (6952 samples); and malicious phishing samples from Nazario Phishing Corpus (9510 samples), Wooyun XSS dataset [WooyunXSSDataset] (168 samples), Trec-07 (50199 samples), and additional 49136 samples generated by summarising and simulating novel types of malicious email attacks that exploit email protocol vulnerabilities. While this dataset is extensive, its heavy reliance on the Enron Corpus for benign emails might introduce bias.
@@
Altwaijry et al. [Altwaijry2024] compared DL models for phishing email detection, using only email body and subject content from Nazario Phishing Corpus and SpamAssassin. They proposed a lightweight 1D-CNN architecture (1D-CNNPD), enhanced with recurrent layers like Bi-GRU. Their best model (1D-CNNPD + Bi-GRU) achieved a 99.66% F1, outperforming traditional ML models and comparable DL approaches such as THEMIS [Fang2019] and DeepAnti-PhishNet [vinayakumar2018deepanti] and demonstrating the effectiveness of CNNs combined with recurrent mechanisms for phishing detection from text. Another recent approach by Koide et al. [Koide2024] introduced ChatSpamDetector, which uses Large Language Models (LLMs), specifically GPT-4, for phishing detection. They prompt GPT-4 directly with raw email text, allowing it to infer phishing characteristics contextually, rather than relying on engineered features. This zero-shot/few-shot approach achieved 99.70% accuracy on a private balanced dataset, showing LLMs’ ability to capture subtle phishing signals.